a father is 5 times as old as his son. In 4 years the sum of their ages will be 56 years. Find their present age
Found 2 solutions by calegebra, richwmiller:
Answer by calegebra(14)

a father is 5 times as old as his son. In 4 years the sum of their ages will be 56 years. Find their present age
let x = son's age
and 5x = father's age
Sum of ages in 4 years = 56
(x + 5x)+ 4 = 56
6x = 52
x = 8.7
5x = 43.3

Answer by richwmiller(17219)

The other tutor forgot to add four to both ages
f=5s
f+4+s+4=56
f+s+8=56
f+s=48
5s+s=48
6s=48
s=8
f=40
check
44+12=56
ok